Two identical conducting spheres, A and B, sit atop insulating stands. When they are touched, 1.63 × 1013 electrons flow from sphere B to sphere A. If the total net charge on the spheres is +3.62 μC, what was the initial charge on sphere B?


1
Expert's answer
2021-03-08T08:26:04-0500

When the sphere A touches the sphere B, "1.63\cdot10^{13}" electrons flow from sphere B to sphere A. Since the charge of the electrons is negative, the sphere A becomes negatively charged:


"q_A=Ne=1.63\cdot10^{13}\cdot(-1.6\cdot10^{-19}\ C)=-2.608\ \mu C."

Then, we can find the initial charge on sphere B:


"q_{tot}=q_A+q_B,""q_B=q_{tot}-q_A=3.62\ \mu C-(-2.608\ \mu C)=6.23\ \mu C."